OpenStudy (anonymous):

can you help me with one other problem : x squared plus y squared plus 14x plus 18y plus 81=0. find the center of the circle.

OpenStudy (anonymous):

(-7,-9)

OpenStudy (anonymous):

Radius 7

OpenStudy (anonymous):

Because (x+7)^2-49=x^2+14x And (y+9)^2-81=y^2+18y So x^2+y^2+14x+18y+81=0 is the same as (x+7)^2+(y+9)^2=7^2

OpenStudy (anonymous):

(x+7)=0 when x=-7 (y+9)=0 when y=-9 center is (-7,-9)

OpenStudy (anonymous):

\[x^2+14x+y^2+18y=-81\] \[(x+7)^2+(y+9)^2=-81+7^2+9^2=-81+49+91=49\] \[(x+7)^2+(y+9)^2=7^2\] formula for circle with center ( h , k ) and radius r is \[(x-h)^2+(y-k)^2=r^2\] and you read off the answer. all the work is "completing the square" \[x^2+14x\] complete by adding 49 to get \[(x+7)^2\] since you add 49 to the left you must add 49 to the right. same for \[y^2+18y\] add 81 to get \[(y+9)^2\]